OpenStudy (anonymous):

The average of 5 numbers is 12. Four of the numbers are 10,13,15 and 21. What is the fifth number

OpenStudy (ash2326):

Let the fifth no. be x , than \[\frac{10+13+15+21+x}{5}=12\] Can you find x now?

OpenStudy (anonymous):

(x+10+13+15+21)/5=12 x+59=12*5 x+59=60 x=1